### v1

# Primal-Dual iLQR

[…]
mm[#1]
Primal-Dual iLQR
João Sousa-Pinto (Apple)
joaospinto@gmail.com
   Dominique Orban (Polytechnique Montréal)
dominique.orban@polymtl.ca
(December 2023)
I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual SQP problem. As our algorithm is a specialization of NPSQP (Gill et al. 1992), it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS (Giftthaler et al. 2018) and FDDP (Mastalli et al. 2020).
II Introduction
II-A Unconstrained Discrete-Time Optimal Control Problems
Unconstrained discrete-time optimal control problems are optimization problems of the form
$&\min_{x_{0},u_{0}\ldots,x_{N}}&\sum\limits_{i=0}^{N-1}g_{i}(x_{i},u_{i})+g_{N}(x_{N})\\ &\mbox{s.t.}&x_{0}=s_{0}\\ &&x_{i+1}=f_{i}(x_{i},u_{i}),\forall i\in\{0,\ldots,N-1\}.\end{aligned}$
Such optimization problems are ubiquitous in the fields of motion planning and controls.
II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP (Mayne 1966), achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in (Shoemaker and Liao 1991). iLQR (Todorov and Li 2004) can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton (Dunn and Bertsekas 1989) exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
DDP, iLQR, and Stagewise Newton are all single shooting methods. This makes them hard to warm start, as the state trajectory cannot be chosen independently of the control trajectory. This can result in more iterations being required for these methods to converge, as well as higher likelihoods of convergence to undesired local optima.
(Giftthaler et al. 2018) introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. (Mastalli et al. 2020) introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
Much work has also been done in handling discrete-time optimal control problems with stagewise constraints. ALTRO (Howell, Jackson, and Manchester 2019), FATROP (Vanroye et al. 2023), and acados (Verschueren et al. 2022) are likely the most competitive open source software packages for this application at the time of writing. ALTRO handles constraints via the augmented Lagrangian method, while FATROP uses an interior point method and acados uses an SQP approach. ForcesNLP (Zanelli et al.) is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
(Singh, Slotine, and Sindhwani 2022) introduces a family of methods combining SQP and DDP, one of which using a similar approach to the one discussed in this paper (what the authors call the “open loop (OL)” variant), which seems to perform poorly in their benchmarks. At several points in their paper and code (available in the Trajax repository on GitHub), there are missing affine terms, for example in the linearized dynamics, which would likely explain the benchmark results.
At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT (Wächter and Biegler 2006) with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$. Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in (Rao, Wright, and Rawlings 1998). The modified KKT system will only have the extra stagewise affine constraints when the problem being solved has other stagewise equality constraints besides the dynamics. A similar algorithm that does not handle other general equality constraints had previously been presented in (Pavlov, Shames, and Manzie 2020). As compared to the specialization of FATROP to problems without stagewise equality constraints, our method has the advantage of not requiring second order corrections for fast local convergence. ForcesNLP also follows an interior point approach, but relies on a Cholesky factorizations instead of LQR decompositions for solving its linear systems.
On the other hand, the augmented Lagrangian technique employed by ALTRO yields linear systems that can be solved via the LQR algorithm, even in the presence of stagewise equality constraints. However, ALTRO has a slightly unorthodox way of doing multiple shooting: it modifies the dynamics $x_{n+1}=f_{n}(x_{n},u_{n})$ into $x_{n+1}=f_{n}(x_{n},u_{n})+e_{n}$ and adds the constraints $e_{n}=0$. This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting method. This approach has the disadvantage of substantially expanding the control dimension.
Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM (Frison and Diehl 2020), an efficient interior point QP solver that exploits the discrete-time optimal control problem structure. While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to (Vanroye et al. 2023). Moreover, due to not employing a globalization strategy, global convergence is not guaranteed. When the functions defining the problem are expensive to evaluate, SQP methods become time-competitive, as the higher number of linear system solves will be compensated by the lower number of evaluations of the nonlinear problem.
II-C Contributions
This paper presents a self-contained description of the specialization of NPSQP to unconstrained discrete-time optimal control problems. As compared to GNMS and FDDP, our method has the advantages of being known to be globally convergent and to not impede superlinear local convergence. Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
This framework can be extended to also handle constrained problems, using augmented Lagrangian (Howell, Jackson, and Manchester 2019), interior point (Vanroye et al. 2023), or SQP (Verschueren et al. 2022) methods. ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses). Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
III General Optimization Background
III-A First-order Optimality Conditions and KKT Systems
Given a constrained optimization problem of the form
$\min_{x}$ $g(x)$ (1)
s.t. $c(x)=0,$ (2)
for any local optimum $x$ meeting the standard linear independence constraint qualification (LICQ) constraints (i.e., for which the rows of $J(c)(x)$ are linearly independent), there exists a vector $\lambda$ (called the Lagrange/KKT multiplier) for which
$\begin{cases}\nabla g(x)+\nabla c(x)\lambda&=0\\ c(x)&=0.\end{cases}$ (3)
In other words, $(x,\lambda)$ is a critical point (but not necessarily a minimizer) of the Lagrangian function $\mathcal{L}(x,\lambda)$ defined as $g(x)+\lambda^{T}c(x)$.
III-B Sequential Quadratic Programming
The sequential quadratic programming (SQP) method for equality-constrained non-linear optimization problems consists of applying Newton’s method for finding zeros of $\nabla\mathcal{L}$, typically in combination with a line search mechanism, in order to ensure global convergence.
Below, let $A=J(c)(x)$, $l=\nabla_{x}\mathcal{L}(x,\lambda)$, and $d=c(x)$. Moreover, let $Q$ be a positive definite approximation of $\nabla_{xx}^{2}\mathcal{L}(x,\lambda)$. Such an approximation can be constructed, for example, via regularization. Alternatively, the terms involving $\nabla_{xx}^{2}c(x)$ can be dropped entirely, and $Q$ would consist purely of a positive definite approximation of $\nabla_{xx}g(x)$. The latter approach, of course, would result in local quadratic convergence being lost. For simplicity of notation, we henceforth drop the dependencies of these terms in $x$.
Each Newton step for finding a zero of $h(z)$ is obtained by solving $h^{\prime}(z)\Delta z=-h(z)$. If $h=\nabla\mathcal{L}$, this corresponds to solving
$\begin{pmatrix}Q&A^{T}\\ A&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}l\\ d\end{pmatrix}.$ (10)
It can be shown (by a simple application of the first-order optimality conditions mentioned above) that $\Delta x$ is the solution of
$\min_{p}$ $\frac{1}{2}p^{T}Qp+l^{T}p$ (11)
s.t. $Ap+d=0.$ (12)
Similarly, it can also be shown that $\Delta\lambda$ is the corresponding Lagrange multiplier. As long as $Q$ is positive definite, this cost function is bounded from below. Depending on the sparsity pattern of this system, weaker requirements on $Q$ might also guarantee this.
Convergence is established when $\Delta x$ and $\Delta\lambda$ are sufficiently close to 00. The solution is feasible as long as $d=0$, but convergence to locally infeasible solutions is possible.
III-C Merit Functions
Merit functions offer a measure of progress towards the final solution and are typically used in combination with a line search procedure. Note that it would be inadequate to simply apply line search on the original objective, as SQP may (and often does) provide a search direction that locally increases that objective (usually trading that off for a reduction in the constraint violations). Filter methods, used in (Wächter and Biegler 2006), provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
Given a penalty parameter $\rho>0$, we define the $\ell_{2}$ merit function
$m_{\rho}(x,\lambda)=\mathcal{L}(x,\lambda)+\frac{\rho}{2}\lVert c(x)\rVert^{2}.$ (13)
Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above. Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&0\end{pmatrix})$ (14)
$=\Delta x^{T}\nabla_{x}m_{\rho}(x,\lambda)$
$=\Delta x^{T}\nabla_{x}\mathcal{L}(x,\lambda)+\rho\Delta x^{T}J(c)(x)^{T}c(x)$
$=-\Delta x^{T}Q\Delta x-\Delta x^{T}A^{T}\Delta\lambda+\rho\Delta x^{T}A^{T}d$
$=-\Delta x^{T}Q\Delta x+\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}\Delta\lambda-\rho\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}d$
$=-\Delta x^{T}Q\Delta x+d^{T}\Delta\lambda-\rho\lVert d\rVert^{2}$
and that
$D(m_{\rho};\begin{pmatrix}0&\Delta\lambda^{T}\end{pmatrix})$ $=\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}\Delta\lambda=d^{T}\Delta\lambda.$ (15)
Therefore:
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})$ (16)
$=$ $-\Delta x^{T}Q\Delta x+2d^{T}\Delta\lambda-\rho\lVert d\rVert^{2}.$
Note that, as $Q$ is positive definite, $-\Delta x^{T}Q\Delta x<0$ unless $\Delta x=0$. When $d$ is not acceptably close to 00, we suggest taking
$\rho=\frac{2\lVert\Delta\lambda\rVert}{\lVert d\rVert},$ (17)
which ensures that $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0$ (due to the Cauchy-Schwarz inequality). Otherwise, we suggest taking $\rho=0.01$.
Note that (Gill et al. 1992) suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0.$ (19)
III-D Line Search
Having computed $\Delta x$, $\Delta\lambda$, and $\rho$, we wish to compute a step size $\alpha$ that results in an acceptable decrease in our merit function $m_{\rho}$. Typically, this is achieved by performing a backtracking line search. $\alpha$ starts as $1$ and decreases by a constant multiplicative factor (often $0.5$) every iteration, until
$m_{\rho}(x+\alpha\Delta x,\lambda+\alpha\Delta\lambda)<$ (20)
$m_{\rho}(x,\lambda)+k\alpha D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})$
is satisfied (where $k\in(0,1)$ is called the Armijo factor; typically, $10^{-4}$ is used). This is called the Armijo condition. This process is guaranteed to terminate as long as $g(x)$ and $c(x)$ are differentiable at $x$ and $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0$.
When $\Delta x=0$ and $d=0$ but $\Delta\lambda\neq 0$, we take a full step without conducting a line search.
Note that NPSQP (Gill et al. 1992) uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
IV Algorithm Derivation
In this section, we specialize the methods described above to the case of unconstrained discrete-time optimal control problems as needed.
We start by defining
$d=\begin{pmatrix}d_{0}\\ \cdots\\ d_{N}\end{pmatrix}=c(x)=\begin{pmatrix}s_{0}-x_{0}\\ f(x_{0},u_{0})-x_{1}\\ \cdots\\ f(x_{N-1},u_{N-1})-x_{N}\end{pmatrix}.$ (28)
The corresponding Lagrange/KKT multipliers are $\lambda_{0},\ldots,\lambda_{N}$ respectively. Letting $x=(x_{0},u_{0},\ldots,x_{N-1},u_{N-1},x_{N})$, $A_{i}=J_{x}(f_{i})(x_{i},u_{i})$ and $B_{i}=J_{u}(f_{i})(x_{i},u_{i})$,
$A=J(c)(x)=$ (29)
$\begin{pmatrix}-I&0&0&0&0&\cdots&0&0&0\\ A_{0}&B_{0}&-I&0&0&\cdots&0&0&0\\ 0&0&A_{1}&B_{1}&-I&\cdots&0&0&0\\ 0&0&0&0&\ddots&\ddots&\ddots&0&0\\ 0&0&0&0&0&\cdots&A_{N-1}&B_{N-1}&-I\end{pmatrix}.$
Note that the LICQ conditions are always met for matrices of this form, due to the presence of the $-I$ blocks, independently of the $A_{i}$ and $B_{i}$.
The SQP constraint $Ap+d=0$ becomes
$\begin{cases}\Delta x_{0}=d_{0}\\ \Delta x_{i+1}=A_{i}\Delta x_{i}+B_{i}\Delta u_{i}+d_{i},\forall i\in\{0,\ldots,N-1\}.\end{cases}$ (30)
Note that if $x_{0}$ is warm started as $s_{0}$ then $\Delta x_{0}$ will always be 00.
Similarly, noting that
$\mathcal{L}(x,\lambda)=$ $\lambda_{0}^{T}(s_{0}-x_{0})+\sum\limits_{i=0}^{N-1}g_{i}(x_{i},u_{i})+$ (31)
$\lambda_{i+1}^{T}(f_{i}(x_{i},u_{i})-x_{i+1})+g_{N}(x_{N}),$
it holds that
$l=\nabla_{x}\mathcal{L}(x,\lambda)=\begin{pmatrix}l_{0}\\ \vdots\\ l_{N}\end{pmatrix}$ (35)
where
$\begin{cases}l_{i}=\nabla g_{i}(x_{i},u_{i})+\nabla f_{i}(x_{i},u_{i})\lambda_{i+1}-\lambda_{i},\\ \forall i\in\{0,\ldots,N-1\}\\ l_{N}=\nabla g_{N}(x_{N})-\lambda_{N}\end{cases}$ (36)
and that
$Q=\nabla_{xx}\mathcal{L}(x,\lambda)=\begin{pmatrix}Q_{0}&&\\ &\ddots\\ &&Q_{N}\end{pmatrix},$ (40)
where
$\begin{cases}Q_{i}=\nabla^{2}g_{i}(x_{i},u_{i})+\nabla^{2}f_{i}(x_{i},u_{i})\lambda_{i+1},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}=\nabla^{2}g_{N}(x_{N}).\end{cases}$ (41)
The SQP cost $\frac{1}{2}p^{T}Qp+l^{T}p$ is simply $\sum\limits_{i=0}^{N}\frac{1}{2}p_{i}^{T}Q_{i}p_{i}+l_{i}^{T}p_{i}$, where $p_{i}=(\Delta x_{i},\Delta u_{i})$ for $i\in\{0,\ldots,N-1\}$ and $p_{N}=\Delta x_{N}$.
If exact Hessians are used, the matrices $Q_{i}$ need not be positive semi-definite. However, we require that positive definite approximations be used instead. It would also be acceptable to only require positive semi-definiteness, as long as the $\nabla_{u_{i}u_{i}}^{2}$ components of the $Q_{i}$ be positive definite. However, this would require checking that $\Delta x^{T}Q\Delta x>0$ at the end of the LQR solve. If this condition is not met, positive definite approximations would have to be used. Requiring positive definiteness in the first place avoids this complication.
In this case, the primal part of the SQP problem is an LQR problem, which can be efficiently solved.
In order to ensure that the matrices $Q_{i}$ are positive semi-definite, it may be helpful to maintain regularization parameters $\mu_{i}$ and always use $Q_{i}+\mu_{i}I$ instead of $Q_{i}$. If, during the LQR process, we detect that $Q_{i}+\mu_{i}I$ is not positive definite, $\mu_{i}$ can be updated by multiplying it by an updated factor $r>1$. When increasing $\mu_{i}$ is not required, we can instead update $\mu_{i}$ by dividing it by $r$. Minimum and maximum regularization parameters may also be established, to allow, when possible, $\mu_{i}$ to eventually be set to 00 if $Q_{i}$ is consistently positive definite, as well as to prevent exploring unreasonably high regularization parameters.
Next, we show how to efficiently compute the $\Delta\lambda$, i.e. how to solve the dual part of the SQP problem. Having determined $\Delta x$, this can be done by specializing $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$. Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\Delta\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
Letting $l_{i}^{x},l_{i}^{u},Q_{i}^{xx},Q_{i}^{xu},Q_{i}^{ux},Q_{i}^{uu}$ be such that
$l_{i}=\begin{pmatrix}l_{i}^{x}\\ l_{i}^{u}\end{pmatrix},Q_{i}=\begin{pmatrix}Q_{i}^{xx}&Q_{i}^{xu}\\ Q_{i}^{ux}&Q_{i}^{uu}\end{pmatrix},\forall i\in\{0,\ldots,N-1\},$ (46)
it holds that
$\begin{cases}Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}-\Delta\lambda_{i}+A_{i}^{T}\Delta\lambda_{i+1}=\\ -l_{i}^{x},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}\Delta x_{N}-\Delta\lambda_{N}=-l_{N}\end{cases}\Rightarrow$ (47)
$\begin{cases}\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x},\\ \forall i\in\{0,\ldots,N-1\}\\ \Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}.\end{cases}$
Note that these equations can be used to recursively compute the $\Delta\lambda_{i}$, in decreasing order of $i$. We call this method the dual LQR backward pass.
Moreover, note that these $\Delta\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\Delta\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
Using the methods described above, we can efficiently compute both $\Delta x$ and $\Delta\lambda$. The remaining parts of the optimization algorithm we describe here don’t require further specialization in the case of unconstrained discrete-time optimal control problems.
V Algorithm Overview
In this section, we provide some pseudocode for the proposed algorithm. For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
def build_lqr(problem, x, u, v):
    n = problem.num_states()
    m = problem.num_controls()
    N = problem.num_stages()
    for i in range(N):
        cost_grad = problem.cost(i)
            .gradient().eval(x[:,i], u[:,i])
        dyn_grad = problem.dynamics(i)
            .gradient().eval(x[:,i], u[:,i])
        A_i = dyn_grad.transpose()[:,:n]
        B_i = dyn_grad.transpose()[:,n:]
        if i == 0:
            d_i = problem.initial_state()
                - x[:,0]
        else:
            y = problem.dynamics(i-1)
                .eval(x[:,i-1], u[:,i-1])
            d_i =  y - x[:,i]
        l_i = cost_grad +
            dyn_grad * costates[:,i+1]
            - costates[:,i]
        cost_hess = problem.cost(i)
            .hessian().eval(x[:,i], u[:,i])
        dyn_hess = problem.dynamics(i)
            .hessians().eval(x[:,i], u[:,i])
        Q_i = cost_hess
            + dyn_hess * costates[:,i+1]
    l_N = problem.cost(N).gradient()
        .eval(x[:,N]) - costates[:,N]
    Q_N = problem.cost(N).hessian()
        .eval(x[:,N])
    return LQR(A_*, B_*, d_*, Q_*, l_*)
def lqr_dual_solve(lqr, dx, du):
    N = lqr.num_stages()
    dv[:,N] = Qxx * dx[:,N] + lx
    for i in reversed(range(N)):
        Qxx = lqr.Q(i).xx()
        Qxu = lqr.Q(i).xu()
        At = lqr.A(i).transpose()
        lx = lqr.l(i).x()
        dv[:,i] = Qxx * dx[:,i] + Qxu * du[:,i]
            + At * dv[:,i+1] + lx
    return dv
def merit_rho(d, dv):
    if d == 0:
        return 0.01
    return 2 * dv.norm() / d.norm()
def eval_merit(problem, x, u, v, rho):
    g = problem.total_cost().eval(x, u)
    d = problem.total_constraints().eval(x, u)
    return g + v.transpose() * d
        + 0.5 * rho * d.squared_norm()
def slope(lqr, dx, du, d, dv, rho):
    N = lqr.num_stages()
    slope = -rho * d.squared_norm()
    for i in range(N):
        dxu = stack(dx[:,i], du[:,i])
        slope += lqr.q(i).transpose() * dxu
        slope += dv[:, i].transpose() * lqr.d(i)
    slope += lqr.q(N).x().transpose() * dx[:, N]
    slope += dv[:, N].transpose() * lqr.d(N)
    return slope
def line_search(problem, lqr, x, u, v,
                dx, du, dv, rho):
    alpha = 1
    m_o = eval_merit(problem, x, u, v, rho)
    slope = slope(lqr, dx, d, dv, rho)
    while alpha > kMinStepSize:
        x_n = x + alpha * dx
        u_n = u + alpha * du
        v_n = v + alpha * dv
        m_n = eval_merit(
            problem, x_n, u_n, v_n, rho)
        if m_n < m_o + kArmijo * alpha * slope:
            return alpha
    raise LineSearchFailure()
def primal_dual_ilqr(problem, states_ws,
        controls_ws, costates_ws):
    x = states_ws
    u = controls_ws
    v = costates_ws
    for it in range(kMaxIterations):
        lqr = build_lqr(problem, x, u)
        dx, du = lqr.Solve()
        dv = lqr_dual_solve(lqr, dx, du)
        if dx == 0 and du == 0:
            if dv == 0:
                return x, u, v
            else:
                v += dv
                continue
        rho = merit_rho(lqr.d(), dv)
        alpha = line_search(
            problem, lqr, x, u,
            v, dx, du, dv, rho)
        x += alpha * dx
        u += alpha * du
        v += alpha * dv
    return x, u, v
VI Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step.
Iteration Objective $\|c\|^{2}$ $m_{\rho}^{\prime}$ $\alpha$
1 69.1912 9.173966407775879 -196341.22 1.0
2 38.890675 10.204395294189453 -282.203 0.5
3 21.515545 9.039937973022461 -99.97078 0.25
4 12.537336 3.5910232067108154 -37.347027 0.5
5 13.005388 1.256274938583374 -7.008366 0.5
6 11.193476 0.05883491784334183 -8.02231 1.0
7 10.798406 0.013367442414164543 -0.79570323 1.0
8 10.654287 0.0059180837124586105 -0.2289545 1.0
9 10.586153 0.001393172424286604 -0.10068051 1.0
10 10.561506 0.0004326050984673202 -0.03127028 1.0
11 10.545758 0.00018692569574341178 -0.01689804 1.0
12 10.535523 8.762026118347421e-05 -0.01067685 1.0
13 10.528571 4.414142313180491e-05 -0.00720222 1.0
14 10.523667 2.3347358364844695e-05 -0.00506186 1.0
15 10.520094 1.2880265785497613e-05 -0.00367544 1.0
16 10.517421 7.388789981632726e-06 -0.0027444 1.0
17 10.515375 4.351216375653166e-06 -0.00209233 1.0
18 10.513782 2.6609566248225747e-06 -0.00163243 1.0
19 10.512507 1.6749986571085174e-06 -0.00129717 1.0
20 10.511484 1.07657911030401e-06 -0.00104529 1.0
21 10.510648 7.084668709467223e-07 -0.00085156 1.0
22 10.509963 4.725115729797835e-07 -0.00069964 1.0
23 10.509394 3.217612913886114e-07 -0.00057881 1.0
24 10.50892 2.2059847992750292e-07 -0.00048149 1.0
25 10.508526 1.53070544683942e-07 -0.00040247 1.0
26 10.508192 1.0723919530164494e-07 -0.00033782 1.0
27 10.507912 7.579520655554006e-08 -0.00028446 1.0
28 10.507676 5.401408387228912e-08 -0.0002403 1.0
29 10.507476 3.848523988381203e-08 -0.00020355 1.0
30 10.507305 2.7843480410183474e-08 -0.00017285 1.0
31 10.50716 2.0009650825159042e-08 -0.00014714 1.0
32 10.507039 1.4696987626905411e-08 -0.00012551 1.0
33 10.506933 1.0591413968086272e-08 -0.00010727 1.0
VII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm. This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in (Gill et al. 1992).
Our algorithm can also be used to significantly speed up ALTRO, one of the state of the art open source software packages for solving constrained discrete-time optimal control problems, by removing the control dimension expansion occurring therein. Our method also removes the need to handle second order corrections, which would simplify and potentially speed up FATROP.

### v2

@@
 We use the LQR algorithm to efficiently solve the primal-dual SQP problem.
-As our algorithm is a specialization of NPSQP (Gill et al.
-1992), it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS (Giftthaler et al.
-2018) and FDDP (Mastalli et al.
-2020).
+As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS [3] and FDDP [6].
 II Introduction
@@
 Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
-The first algorithm to improve on this was DDP (Mayne 1966), achieving complexity $O(N(n+m)^{3})$.
+The first algorithm to improve on this was DDP [7], achieving complexity $O(N(n+m)^{3})$.
 Its convergence properties were studied in (Shoemaker and Liao 1991).
-iLQR (Todorov and Li 2004) can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation.
-Stagewise Newton (Dunn and Bertsekas 1989) exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
+iLQR [11] can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation.
+Stagewise Newton [1] exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
 DDP, iLQR, and Stagewise Newton are all single shooting methods.
@@
 This can result in more iterations being required for these methods to converge, as well as higher likelihoods of convergence to undesired local optima.
-(Giftthaler et al.
-2018) introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above.
+Moreover, they require that the dynamics be evaluated sequentially, making them harder to parallelize.
+[3] introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above.
 There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence.
-(Mastalli et al.
-2020) introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes.
+[6] introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes.
 There are also no available convergence results for FDDP.
@@
 Much work has also been done in handling discrete-time optimal control problems with stagewise constraints.
-ALTRO (Howell, Jackson, and Manchester 2019), FATROP (Vanroye et al.
-2023), and acados (Verschueren et al.
-2022) are likely the most competitive open source software packages for this application at the time of writing.
+ALTRO [5], FATROP [12], and acados [13] are likely the most competitive open source software packages for this application at the time of writing.
 ALTRO handles constraints via the augmented Lagrangian method, while FATROP uses an interior point method and acados uses an SQP approach.
-ForcesNLP (Zanelli et al.) is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
-(Singh, Slotine, and Sindhwani 2022) introduces a family of methods combining SQP and DDP, one of which using a similar approach to the one discussed in this paper (what the authors call the “open loop (OL)” variant), which seems to perform poorly in their benchmarks.
-At several points in their paper and code (available in the Trajax repository on GitHub), there are missing affine terms, for example in the linearized dynamics, which would likely explain the benchmark results.
-At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT (Wächter and Biegler 2006) with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$.
-Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in (Rao, Wright, and Rawlings 1998).
+ForcesNLP [15] is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
+At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT [14] with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$.
+Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in [9].
 The modified KKT system will only have the extra stagewise affine constraints when the problem being solved has other stagewise equality constraints besides the dynamics.
-A similar algorithm that does not handle other general equality constraints had previously been presented in (Pavlov, Shames, and Manzie 2020).
+A similar algorithm that does not handle other general equality constraints had previously been presented in [8].
 As compared to the specialization of FATROP to problems without stagewise equality constraints, our method has the advantage of not requiring second order corrections for fast local convergence.
@@
 However, ALTRO has a slightly unorthodox way of doing multiple shooting: it modifies the dynamics $x_{n+1}=f_{n}(x_{n},u_{n})$ into $x_{n+1}=f_{n}(x_{n},u_{n})+e_{n}$ and adds the constraints $e_{n}=0$.
-This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting method.
+This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting algorithm.
 This approach has the disadvantage of substantially expanding the control dimension.
-Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM (Frison and Diehl 2020), an efficient interior point QP solver that exploits the discrete-time optimal control problem structure.
-While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to (Vanroye et al.
-2023).
+Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM [2], an efficient interior point QP solver that exploits the discrete-time optimal control problem structure.
+While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to [12].
 Moreover, due to not employing a globalization strategy, global convergence is not guaranteed.
@@
 Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
-This framework can be extended to also handle constrained problems, using augmented Lagrangian (Howell, Jackson, and Manchester 2019), interior point (Vanroye et al.
-2023), or SQP (Verschueren et al.
-2022) methods.
+This framework can be extended to also handle constrained problems, using augmented Lagrangian [5], interior point [12], or SQP [13] methods.
 ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses).
@@
 Note that it would be inadequate to simply apply line search on the original objective, as SQP may (and often does) provide a search direction that locally increases that objective (usually trading that off for a reduction in the constraint violations).
-Filter methods, used in (Wächter and Biegler 2006), provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
+Filter methods, used in  [14], provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
 Given a penalty parameter $\rho>0$, we define the $\ell_{2}$ merit function
@@
 which ensures that $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0$ (due to the Cauchy-Schwarz inequality). Otherwise, we suggest taking $\rho=0.01$.
-Note that (Gill et al. 1992) suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
+Note that [4] suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
 $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0.$ (19)
@@
 When $\Delta x=0$ and $d=0$ but $\Delta\lambda\neq 0$, we take a full step without conducting a line search.
-Note that NPSQP (Gill et al. 1992) uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
+Note that NPSQP [4] uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
 IV Algorithm Derivation
@@
 We also increased the final state cost gains on the positions to $1000$.
+This seems to be the problem used in [16].
 Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step.
+A JAX implementation of our algorithm takes 46ms to finish solving this problem.
+Below, you can find a visual description of the problem and solution, as well as solver logs.
+Figure 1: The solution of the quad-pendulum problem.
 Iteration Objective $\|c\|^{2}$ $m_{\rho}^{\prime}$ $\alpha$
@@
 This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
-This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in (Gill et al.
-1992).
+This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in [4].
+Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
+Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton.
+Moreover, it allows that the evaluation of the dynamics and derivatives be computed in parallel.
 Our algorithm can also be used to significantly speed up ALTRO, one of the state of the art open source software packages for solving constrained discrete-time optimal control problems, by removing the control dimension expansion occurring therein.

### v3

@@
 Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function.
-We use the LQR algorithm to efficiently solve the primal-dual SQP problem.
-As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS [3] and FDDP [6].
+We use the LQR algorithm to efficiently solve the primal component of the Newton-KKT system, and use a dual LQR backward pass to solve its dual component.
+We also present a new parallel algorithm for solving the dual component of the Newton-KKT system in $O(\log(N))$ parallel time, where $N$ is the number of stages.
+Combining it with [17], we are able to solve the full Newton-KKT system in $O(\log(N))$ parallel time.
+The remaining parts of our method have constant parallel time complexity per iteration.
+Therefore, this paper provides, for the first time, a practical, highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
+As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6].
 II Introduction
@@
 II-C Contributions
-This paper presents a self-contained description of the specialization of NPSQP to unconstrained discrete-time optimal control problems.
-As compared to GNMS and FDDP, our method has the advantages of being known to be globally convergent and to not impede superlinear local convergence.
-Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
-This framework can be extended to also handle constrained problems, using augmented Lagrangian [5], interior point [12], or SQP [13] methods.
-ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses).
-Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
+This paper introduces the Primal-Dual iLQR algorithm, which consists of a specialization of the generic NPSQP algorithm to the case of unconstrained discrete-time optimal control problems.
+Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
+We also introduce a new method for parallelizing the computation of the dual solution of the Newton-KKT systems, with parallel time complexity $O(\log(N)\log(n))$, by framing the dual LQR backward pass as an associative scan [18].
+This can be combined with the method of [17] to solve the primal-dual Newton-KKT system in parallel time complexity logarithmic in $N$.
+Our algorithm improves on earlier direct multiple shooting methods, specifically by guaranteeing global convergence and not impeding local superlinear convergence (even without second order corrections), under the assumptions described in [4].
 III General Optimization Background
@@
 The remaining parts of the optimization algorithm we describe here don’t require further specialization in the case of unconstrained discrete-time optimal control problems.
-V Algorithm Overview
-In this section, we provide some pseudocode for the proposed algorithm.
+V Parallel Dual LQR Backward Pass with Associative Scans
+Associative scans are a common parallelization mechanism used in functional programming, first introduced in [18].
+They were used in [17] to derive a simple method for solving (primal) LQR problems in $O(\log(m)+\log(N)\log(n))$ parallel time, where $N,n,m$ are respectively the number of stages, states, and controls.
+In this section, we give a brief description of what associative scans are, and how they can also be used to derive a similar method that is equivalent to the dual LQR backward pass, while only taking $O(\log(N)\log(n))$ parallel time.
+Note that the $\log(n)$ factor here comes from the parallel time complexity of taking the product of two matrices of dimension $n$.
+Given a set $\mathcal{X}$, a function $f:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ is said to be associative if $\forall a,b,c\in\mathcal{X},f(f(a,b),c)=f(a,f(b,c))$.
+The forward associative scan operation $S_{f}(x_{1},\ldots,x_{n};f)$ can be recursively defined by $S_{f}(x_{1};f)=x_{1}$ and $S_{f}(x_{1},\ldots,x_{i+1};f)=(y_{1},\ldots,y_{i},f(y_{i},x_{i+1}))$, where $y_{1},\ldots,y_{i}=S_{f}(x_{1},\ldots,x_{i};f)$.
+Similarly, the reverse associative scan operation $S_{r}(x_{1},\ldots,x_{n};f)$ can be recursively defined by $S_{r}(x_{1};f)=x_{1}$ and $S_{r}(x_{1},\ldots,x_{i+1};f)=(f(x_{1},y_{2}),y_{2},\ldots,y_{i+1})$, where $y_{2},\ldots,y_{i+1}=S_{r}(x_{2},\ldots,x_{i+1};f)$.
+[18] provides a method for performing associative scans of $N$ elements in parallel time $O(\log(N))$.
+As shown in eq.
+47, the dual solutions $\Delta\lambda_{i}$ of the Newton-KKT system eq.
+10 can be obtained by applying a backward pass starting at $\Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}$, with recursion rule $\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x}$.
+We can simplify the notation by defining, for $i\in\{0,\ldots,N-1\}$, $M_{i}=A_{i}^{T}$ and $m_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+l_{i}^{x}$, so that the recursion rule becomes simply $\Delta\lambda_{i}=M_{i}\Delta\lambda_{i+1}+m_{i}$.
+Moreover, we define $m_{N}=Q_{N}\Delta x_{N}+l_{N}$ and $M_{N}=I$.
+Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(a+Bc,BD)$, we claim that $f$ is associative and that $\Delta\lambda_{0},\ldots,\Delta\lambda_{N}$ can be obtained by doing a reverse associative scan with respect to $f$ on $(m_{0},M_{0}),\ldots,(m_{N-1},M_{N-1}),(m_{N},M_{N})$.
+The associative property is simple to verify:
+$f(f((a,B),(c,D)),(e,F))$ (48)
+$=$ $f((a+Bc,BD),(e,F))$
+$=$ $(a+Bc+BDe,BDF)$
+$=$ $(a+B(c+De),BDF)$
+$=$ $f((a,B),(c+De,DF))$
+$=$ $f((a,B),f((c,D),(e,F))).$
+Let $F$ denote the operator that extracts the first element from its input sequence. We can show by induction that
+$(\Delta\lambda_{i},M_{i}\cdots M_{N})=F(S_{r}((m_{i},M_{i})),\ldots,(m_{N},M_{N});f).$ (49)
+This is clearly true for $i=N$, as
+$S_{r}((m_{N},M_{N});f)=(m_{N},M_{N})=(\Delta\lambda_{N},M_{N}).$ (50)
+For $i<N$, by definition of $S_{r}$ and the induction hypothesis,
+$F(S_{r}((m_{i},M_{i}),\ldots,(m_{N},M_{N});f))$ (51)
+$=$ $f((m_{i},M_{i}),F(S_{r}((m_{i+1},M_{i+1}),\ldots,(m_{N},M_{N});f)))$
+$=$ $f((m_{i},M_{i}),(\Delta\lambda_{i+1},M_{i+1}\cdots M_{N}))$
+$=$ $(m_{i}+M_{i}\Delta\lambda_{i+1},M_{i}\cdots M_{N})$
+$=$ $(\Delta\lambda_{i},M_{i}\cdots M_{N}).$
+VI Algorithm Overview
+In this section, we provide some pseudocode for the sequential version of the proposed algorithm.
 For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
@@
     return x, u, v
-VI Benchmarks
+VII Benchmarks
 Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$).
@@
 Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step.
-A JAX implementation of our algorithm takes 46ms to finish solving this problem.
+A JAX implementation of our algorithm takes 46ms to finish solving this problem (on the CPU of an M2 MacBook Air).
 Below, you can find a visual description of the problem and solution, as well as solver logs.
@@
 33 10.506933 1.0591413968086272e-08 -0.00010727 1.0
-VII Conclusion
+VIII Conclusion
 This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
-This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in [4].
-Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
+Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
+Moreover, it inherits all the desirable properties that NPSQP has been shown to have, such as not impeding local superlinear convergence, global convergence, and the lack of need for second order corrections.
+None of acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6] combine all of these properties.
+The proposed algorithm has sequential time complexity $O(N(n+m)^{3})$, where $N,n,m$ are respectively the number of stages, states, and controls of the given problem, and parallel time complexity $O(\log(m)+\log(N)\log(n)$.
+Note that a naïf application of the NPSQP algorithm would have sequential time complexity $O(N^{3}(n+m)^{3})$ and parallel time complexity $O(\log(N)+\log(n+m))$ (although requiring a much larger number of workers), as the Newton-KKT systems have dimension $(N+1)(n+m)$.
+This is due to the fact that inverting a matrix of dimension $d$ takes $O(d^{3})$ sequential time and $O(\log(d))$ parallel time.
+This framework can easily be extended to also handle constrained problems, ideally using an augmented Lagrangian [5] or interior point [12] method; using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
 Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton.
-Moreover, it allows that the evaluation of the dynamics and derivatives be computed in parallel.
-Our algorithm can also be used to significantly speed up ALTRO, one of the state of the art open source software packages for solving constrained discrete-time optimal control problems, by removing the control dimension expansion occurring therein.
-Our method also removes the need to handle second order corrections, which would simplify and potentially speed up FATROP.
+Moreover, as it never performs dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.

### v4

@@
 Primal-Dual iLQR
-[…]
-mm[#1]
 Primal-Dual iLQR
@@
 Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function.
-We use the LQR algorithm to efficiently solve the primal component of the Newton-KKT system, and use a dual LQR backward pass to solve its dual component.
-We also present a new parallel algorithm for solving the dual component of the Newton-KKT system in $O(\log(N))$ parallel time, where $N$ is the number of stages.
-Combining it with [17], we are able to solve the full Newton-KKT system in $O(\log(N))$ parallel time.
-The remaining parts of our method have constant parallel time complexity per iteration.
-Therefore, this paper provides, for the first time, a practical, highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
+We use the LQR algorithm to efficiently solve the primal-dual Newton-KKT system.
 As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6].
+As our algorithm avoids sequential rollouts of the nonlinear dynamics, it can be combined with [17] to run in $O(\log(N))$ parallel time per iteration (where $N$ is the number of stages), as well as $O(1)$ parallel time per line search iteration.
+Therefore, this paper provides a practical, theoretically sound, and highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
 II Introduction
@@
 The first algorithm to improve on this was DDP [7], achieving complexity $O(N(n+m)^{3})$.
-Its convergence properties were studied in (Shoemaker and Liao 1991).
+Its convergence properties were studied in [10].
 iLQR [11] can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation.
@@
 [3] introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above.
-There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence.
+There are no available convergence results for iLQR-GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence.
 [6] introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes.
@@
 Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
-We also introduce a new method for parallelizing the computation of the dual solution of the Newton-KKT systems, with parallel time complexity $O(\log(N)\log(n))$, by framing the dual LQR backward pass as an associative scan [18].
-This can be combined with the method of [17] to solve the primal-dual Newton-KKT system in parallel time complexity logarithmic in $N$.
+Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n))$ parallel time complexity, using a minor variation of [18], and that the remaining parts of our method have constant parallel time complexity.
 Our algorithm improves on earlier direct multiple shooting methods, specifically by guaranteeing global convergence and not impeding local superlinear convergence (even without second order corrections), under the assumptions described in [4].
@@
 Requiring positive definiteness in the first place avoids this complication.
-In this case, the primal part of the SQP problem is an LQR problem, which can be efficiently solved.
+In this case, the the SQP problem is a primal-dual LQR problem, which can be efficiently solved.
 In order to ensure that the matrices $Q_{i}$ are positive semi-definite, it may be helpful to maintain regularization parameters $\mu_{i}$ and always use $Q_{i}+\mu_{i}I$ instead of $Q_{i}$.
@@
 Minimum and maximum regularization parameters may also be established, to allow, when possible, $\mu_{i}$ to eventually be set to 00 if $Q_{i}$ is consistently positive definite, as well as to prevent exploring unreasonably high regularization parameters.
-Next, we show how to efficiently compute the $\Delta\lambda$, i.e.
-how to solve the dual part of the SQP problem.
-Having determined $\Delta x$, this can be done by specializing $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$.
-Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\Delta\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
-Letting $l_{i}^{x},l_{i}^{u},Q_{i}^{xx},Q_{i}^{xu},Q_{i}^{ux},Q_{i}^{uu}$ be such that
-$l_{i}=\begin{pmatrix}l_{i}^{x}\\ l_{i}^{u}\end{pmatrix},Q_{i}=\begin{pmatrix}Q_{i}^{xx}&Q_{i}^{xu}\\ Q_{i}^{ux}&Q_{i}^{uu}\end{pmatrix},\forall i\in\{0,\ldots,N-1\},$ (46)
-it holds that
-$\begin{cases}Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}-\Delta\lambda_{i}+A_{i}^{T}\Delta\lambda_{i+1}=\\ -l_{i}^{x},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}\Delta x_{N}-\Delta\lambda_{N}=-l_{N}\end{cases}\Rightarrow$ (47)
-$\begin{cases}\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x},\\ \forall i\in\{0,\ldots,N-1\}\\ \Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}.\end{cases}$
-Note that these equations can be used to recursively compute the $\Delta\lambda_{i}$, in decreasing order of $i$. We call this method the dual LQR backward pass.
-Moreover, note that these $\Delta\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\Delta\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
-Using the methods described above, we can efficiently compute both $\Delta x$ and $\Delta\lambda$.
-The remaining parts of the optimization algorithm we describe here don’t require further specialization in the case of unconstrained discrete-time optimal control problems.
-V Parallel Dual LQR Backward Pass with Associative Scans
+Another option would be to regularize the $Q_{i}$ by performing an explicit eigenvalue decomposition and removing negative (or non-positive, when positive definiteness is required) eigenvalues.
+V Sequential Primal LQR Overview
+In this section, we will go over the conventional sequential algorithm for solving primal LQR problems.
+In the interest of not deviating from standard notation, variable names may henceforth not match earlier parts of this paper.
+A linear-quadratic regulator (LQR) problem is an unconstrained discrete-time optimal control problem where the costs are quadratic functions and the dynamics are affine functions.
+Specifically, they are optimization problems of the form
+$\min_{x_{0},u_{0}\ldots,x_{N}}$ $\sum\limits_{i=0}^{N-1}\left(\frac{1}{2}x_{i}^{T}Q_{i}x_{i}+q_{i}^{T}x_{i}+\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}\right.$ (42)
+$+x_{i}^{T}M_{i}u_{i}\right)+\frac{1}{2}x_{N}^{T}Q_{N}x_{N}+q_{N}^{T}x_{N}$
+s.t. $x_{0}=s_{0}$
+$x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i},\forall i\in\{0,\ldots,N-1\}.$
+Note that the $R_{i}$ are required to be positive definite, and that the $Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T}$ are required to be positive semi-definite, otherwise a minimum may not exist.
+The conventional method for solving LQR problems sequentially relies on the fact that the optimal cost-to-go functions are quadratic on the state at the corresponding stage.
+Therefore, such functions $V_{i}(x_{i})=0.5x_{i}^{T}P_{i}x_{i}+p_{i}^{T}x_{i}+z_{i}$ can be computed in decreasing order of stage (i.e.
+$i$).
+Moreover, the optimal controls to be applied at each stage can be shown to be affine functions of the state at the corresponding stage, i.e.
+$u_{i}=K_{i}x_{i}+k_{i}$.
+The $K_{i}$ and $k_{i}$ can be computed as part of the same backward pass as the $P_{i}$ and $p_{i}$.
+Note that the constants $z_{i}$ do not need to be computed.
+Initializing $P_{N}=Q_{N}$ and $p_{N}=q_{N}$, and introducing $G_{i},H_{i},h_{i}$ as convenient auxiliary variables, the following recursion rules can be used:
+$G_{i}=R_{i}+B_{i}^{T}P_{i+1}B_{i}$ (43)
+$H_{i}=B_{i}^{T}+P_{i+1}A_{i}+M_{i}^{T}$
+$h_{i}=B_{i}^{T}(p_{i+1}+P_{i+1}c_{i})+r_{i}$
+$K_{i}=-G_{i}^{-1}H_{i}$
+$k_{i}=-G_{i}^{-1}h_{i}$
+$P_{i}=Q_{i}+A_{i}^{T}P_{i+1}A_{i}+K_{i}^{T}H_{i}$
+$p_{i}=q_{i}+A_{i}^{T}(p_{i+1}+P_{i+1}c_{i})+K_{i}^{T}h_{i}$
+Once the $K_{i},k_{i}$ have been computed, the $u_{i}$ can be computed in increasing order of $i$, by alternating evaluations of $u_{i}=K_{i}x_{i}+k_{i}$ and $x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i}$.
+This is usually called the LQR forward pass.
+VI Associative Scans Overview
 Associative scans are a common parallelization mechanism used in functional programming, first introduced in [18].
 They were used in [17] to derive a simple method for solving (primal) LQR problems in $O(\log(m)+\log(N)\log(n))$ parallel time, where $N,n,m$ are respectively the number of stages, states, and controls.
-In this section, we give a brief description of what associative scans are, and how they can also be used to derive a similar method that is equivalent to the dual LQR backward pass, while only taking $O(\log(N)\log(n))$ parallel time.
-Note that the $\log(n)$ factor here comes from the parallel time complexity of taking the product of two matrices of dimension $n$.
 Given a set $\mathcal{X}$, a function $f:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ is said to be associative if $\forall a,b,c\in\mathcal{X},f(f(a,b),c)=f(a,f(b,c))$.
@@
 [18] provides a method for performing associative scans of $N$ elements in parallel time $O(\log(N))$.
-As shown in eq.
-47, the dual solutions $\Delta\lambda_{i}$ of the Newton-KKT system eq.
-10 can be obtained by applying a backward pass starting at $\Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}$, with recursion rule $\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x}$.
-We can simplify the notation by defining, for $i\in\{0,\ldots,N-1\}$, $M_{i}=A_{i}^{T}$ and $m_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+l_{i}^{x}$, so that the recursion rule becomes simply $\Delta\lambda_{i}=M_{i}\Delta\lambda_{i+1}+m_{i}$.
-Moreover, we define $m_{N}=Q_{N}\Delta x_{N}+l_{N}$ and $M_{N}=I$.
-Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(a+Bc,BD)$, we claim that $f$ is associative and that $\Delta\lambda_{0},\ldots,\Delta\lambda_{N}$ can be obtained by doing a reverse associative scan with respect to $f$ on $(m_{0},M_{0}),\ldots,(m_{N-1},M_{N-1}),(m_{N},M_{N})$.
-The associative property is simple to verify:
+VII Parallel Primal LQR Overview
+A parallel algorithm for solving these problems was presented in [17], although we require some minor changes due to a difference in problem formulations.
+In order to describe this method, we need to introduce an important definition.
+An interval value function $V_{i\rightarrow j}(x_{i},x_{j})$ maps states $x_{i},x_{j}$ at stages $i,j$ to the minimum possible cost incurred in stages $i,\ldots,j-1$ among all trajectories that start at state $x_{i}$ in stage $i$ and end at state $x_{j}$ in stage $j$ ($\infty$ if no such trajectory exists).
+The key insight of [17] is that the optimal interval value functions can be represented as
+$V_{i\rightarrow j}(x_{i},x_{j})=\max\limits_{\lambda}\left(\frac{1}{2}x_{i}^{T}P_{i\rightarrow j}x_{i}+p_{i\rightarrow j}^{T}x_{i}\right.$ (44)
+$\left.-\frac{1}{2}\lambda^{T}C_{i\rightarrow j}\lambda-\lambda^{T}\left(x_{j}-A_{i\rightarrow j}x_{i}-c_{i\rightarrow j}\right)\right).$
+Initializing, for $i\in\{0,\ldots,N-1\}$, $V_{i\rightarrow i+1}$ with
+$P_{i\rightarrow i+1}$ $=Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T},$ (45)
+$p_{i\rightarrow i+1}$ $=q_{i}-M_{i}R_{i}^{-1}r_{i},$
+$A_{i\rightarrow i+1}$ $=A_{i}-B_{i}R_{i}^{-1}M_{i}^{T},$
+$C_{i\rightarrow i+1}$ $=B_{i}R_{i}^{-1}B_{i}^{T},$
+$c_{i\rightarrow i+1}$ $=c_{i}-B_{i}R_{i}^{-1}r_{i},$
+and initializing $V_{N\rightarrow N+1}$ with
+$P_{N\rightarrow N+1}$ $=Q_{N},$ (46)
+$p_{N\rightarrow N+1}$ $=q_{N},$
+$A_{N\rightarrow N+1}$ $=0,$
+$C_{N\rightarrow N+1}$ $=0,$
+$c_{N\rightarrow N+1}$ $=0,$
+the following combination rules can be applied to compute $V_{i\rightarrow k}$ from $V_{i\rightarrow j}$ and $V_{j\rightarrow k}$:
+$P_{i\rightarrow k}$ $=A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}P_{j\rightarrow k}A_{i\rightarrow j}+P_{i\rightarrow j},$ (47)
+$p_{i\rightarrow k}$ $=A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}\left(p_{j\rightarrow k}+J_{j\rightarrow k}c_{i\rightarrow j}\right)$
+$+p_{i\rightarrow j},$
+$A_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}A_{i\rightarrow j},$
+$C_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}C_{i\rightarrow j}A_{j\rightarrow k}^{T}+C_{j\rightarrow k},$
+$c_{i\rightarrow k}$ $=A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}\left(c_{i\rightarrow j}-C_{i\rightarrow j}p_{j\rightarrow k}\right)$
+$+c_{j\rightarrow k}.$
+A reverse associative scan [18] can be used to compute the $P_{i\rightarrow N+1},p_{i\rightarrow N+1}$ (i.e.
+the $P_{i},p_{i}$ from section V), which in turn can be used to compute the $K_{i},k_{i}$ in $O(1)$ parallel time.
+Finally, we wish to compute the $x_{i},u_{i}$ from the $K_{i},k_{i}$.
+Note that the sequential LQR forward pass has $O(T)$ parallel time complexity.
+However, as done in [17], we can reduce the computation of the $x_{i}$ to a sequential composition of affine functions, which can also be parallelized with an associative scan [18].
+This will be described in the next section.
+$x_{i+1}$ $=A_{i}x_{i}+B_{i}u_{i}+c_{i}=A_{i}x_{i}+B_{i}\left(K_{i}x_{i}+k_{i}\right)+c_{i}$ (48)
+$=\left(A_{i}+B_{i}K_{i}\right)x_{i}+\left(B_{i}k_{i}+c_{i}\right)$
+Once these affine functions have been composed, they can be independently applied to $x_{0}$ to recover all the $x_{i}$.
+The $u_{i}$ can then be computed in $O(1)$ parallel time by independently evaluating $u_{i}=K_{i}x_{i}+k_{i}$.
+VIII Composing Affine Functions with Associative Scans
+In this section, we describe an $O(\log(N))$ parallel time algorithm for composing $N$ affine functions $F_{i}(x)=M_{i}x_{i}+m_{i}$, as done in [17].
+Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(Da+c,DB)$, we claim that $f$ is associative.
+This is simple to verify:
 $f(f((a,B),(c,D)),(e,F))$ (49)
-$=$ $f((a+Bc,BD),(e,F))$
-$=$ $(a+Bc+BDe,BDF)$
-$=$ $(a+B(c+De),BDF)$
-$=$ $f((a,B),(c+De,DF))$
+$=$ $f((Da+c,DB),(e,F))$
+$=$ $(F(Da+c)+e,F(DB))$
+$=$ $((FD)a+Fc+e,(FD)B)$
+$=$ $f((a,B),(Fc+e,FD))$
 $=$ $f((a,B),f((c,D),(e,F))).$
-Let $F$ denote the operator that extracts the first element from its input sequence. We can show by induction that
-$(\Delta\lambda_{i},M_{i}\cdots M_{N})=F(S_{r}((m_{i},M_{i})),\ldots,(m_{N},M_{N});f).$ (49)
-This is clearly true for $i=N$, as
-$S_{r}((m_{N},M_{N});f)=(m_{N},M_{N})=(\Delta\lambda_{N},M_{N}).$ (50)
-For $i<N$, by definition of $S_{r}$ and the induction hypothesis,
-$F(S_{r}((m_{i},M_{i}),\ldots,(m_{N},M_{N});f))$ (51)
-$=$ $f((m_{i},M_{i}),F(S_{r}((m_{i+1},M_{i+1}),\ldots,(m_{N},M_{N});f)))$
-$=$ $f((m_{i},M_{i}),(\Delta\lambda_{i+1},M_{i+1}\cdots M_{N}))$
-$=$ $(m_{i}+M_{i}\Delta\lambda_{i+1},M_{i}\cdots M_{N})$
-$=$ $(\Delta\lambda_{i},M_{i}\cdots M_{N}).$
-VI Algorithm Overview
+Moreover, note that $f$ is the affine function composition operator, as $D(Bx+a)+c=(DB)x+(Da+c)$.
+IX Dual LQR Backward Pass
+Next, we show how to compute the multipliers $\lambda$, i.e.
+how to solve the dual part of the LQR problem.
+Note that primal-dual LQR problems can be solved in a single Newton step, independently of the starting values of $x,\lambda$.
+Therefore, starting with $x,\lambda=0$, the $\Delta x,\Delta\lambda$ from the Newton-KKT system correspond to the optimal $x,\lambda$.
+Having determined $x$ using one of the primal LQR algorithms, determining $\lambda$ can be done by specializing $\nabla_{xx}\mathcal{L}(0)x+J(c)(0)^{T}\lambda=-\nabla_{x}\mathcal{L}(0)$ (i.e.
+the first block-row of the Newton-KKT system).
+Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
+Specifically, for $i\in\{0,\ldots,N-1\}$, it holds that
+$Q_{i}x_{i}+M_{i}u_{i}-\lambda_{i}+A_{i}^{T}\lambda_{i+1}=-q_{i}\Rightarrow$ (50)
+$\lambda_{i}=Q_{i}x_{i}+M_{i}u_{i}+A_{i}^{T}\lambda_{i+1}+q_{i},$
+and that
+$Q_{N}x_{N}-\lambda_{N}=-q_{N}\Rightarrow$ (51)
+$\lambda_{N}=Q_{N}x_{N}+q_{N}.$
+Note that these equations can be used to recursively compute the $\lambda_{i}$, in decreasing order of $i$.
+Moreover, note that these $\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(0)x+J(c)(0)^{T}\lambda=-\nabla_{x}\mathcal{L}(0)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
+X Parallel Dual LQR
+While the dual LQR backward pass described in section IX can be parallelized with a reverse associative scan (in a similar fashion to section VIII), we can actually compute the multipliers $\lambda$ in $O(1)$ parallel time.
+The multiplier $\lambda_{0}$ is also the Lagrange/KKT multiplier of
+$\min_{x_{0}}\frac{1}{2}x_{0}^{T}P_{0}x_{0}+p_{0}^{T}x_{0}\quad\mbox{s.t.}\quad x_{0}=s_{0}.$ (52)
+By the first order optimality conditions, the gradient of the corresponding Lagrangian
+$\mathcal{L}(x_{0},\lambda_{0})=\frac{1}{2}x_{0}^{T}P_{0}x_{0}+p_{0}^{T}x_{0}+\lambda^{T}(s_{0}-x_{0})$ (53)
+must be 00. In particular,
+$\nabla_{x_{0}}\mathcal{L}(x_{0},\lambda_{0})=0\Rightarrow\lambda_{0}=P_{0}x_{0}+p_{0}.$ (54)
+Similarly, each multiplier $\lambda_{i}$ is also the Lagrange/KKT multiplier of
+$\min_{u_{i},x_{i+1}}$ $\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}+x_{i}^{T}M_{i}u_{i}$ (55)
+$+\frac{1}{2}x_{i+1}^{T}P_{i+1}x_{i+1}+p_{i+1}^{T}x_{i+1}$
+s.t. $x_{i+1}=A_{i}x_{i}+B_{i}u_{i}+c_{i}.$
+The corresponding Lagrangian is
+$\mathcal{L}(u_{i},x_{i+1},\lambda_{i+1})$ $=\frac{1}{2}u_{i}^{T}R_{i}u_{i}+r_{i}^{T}u_{i}+x_{i}^{T}M_{i}u_{i}$ (56)
+$+\frac{1}{2}x_{i+1}^{T}P_{i+1}x_{i+1}+p_{i+1}^{T}x_{i+1}$
+$+\lambda_{i+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{i}-x_{i+1}\right).$
+Similarly, the first order optimality conditions imply
+$\nabla_{x_{i+1}}\mathcal{L}(u_{i},x_{i+1},\lambda_{i+1})=0\Rightarrow\lambda_{i+1}=P_{i+1}x_{i+1}+p_{i+1}.$ (57)
+Finally, note that all the $\lambda_{i}$ can be evaluated independently, so a backward pass is not required using these formulas.
+This method of computing the multipliers also has the advantage of involving fewer computations and being more numerically stable.
+XI Algorithm Overview
 In this section, we provide some pseudocode for the sequential version of the proposed algorithm.
@@
         cost_grad = problem.cost(i)
-            .gradient().eval(x[:,i], u[:,i])
+            .gradient()
+            .eval(x[:,i], u[:,i])
         dyn_grad = problem.dynamics(i)
-            .gradient().eval(x[:,i], u[:,i])
-        A_i = dyn_grad.transpose()[:,:n]
-        B_i = dyn_grad.transpose()[:,n:]
+            .gradient()
+            .eval(x[:,i], u[:,i])
+        A_i = dyn_grad.T[:,:n]
+        B_i = dyn_grad.T[:,n:]
         if i == 0:
@@
         cost_hess = problem.cost(i)
-            .hessian().eval(x[:,i], u[:,i])
+            .hessian()
+            .eval(x[:,i], u[:,i])
         dyn_hess = problem.dynamics(i)
-            .hessians().eval(x[:,i], u[:,i])
+            .hessians()
+            .eval(x[:,i], u[:,i])
         Q_i = cost_hess
@@
     return LQR(A_*, B_*, d_*, Q_*, l_*)
-def lqr_dual_solve(lqr, dx, du):
-    N = lqr.num_stages()
-    dv[:,N] = Qxx * dx[:,N] + lx
-    for i in reversed(range(N)):
-        Qxx = lqr.Q(i).xx()
-        Qxu = lqr.Q(i).xu()
-        At = lqr.A(i).transpose()
-        lx = lqr.l(i).x()
-        dv[:,i] = Qxx * dx[:,i] + Qxu * du[:,i]
-            + At * dv[:,i+1] + lx
-    return dv
+def lqr_dual_solve(P, p, dx):
+    Np1 = dx.shape[1]
+    return [P[i] * dx[:, i] + p[i]
+        for i in range(Np1)]
 def merit_rho(d, dv):
@@
     g = problem.total_cost().eval(x, u)
-    d = problem.total_constraints().eval(x, u)
-    return g + v.transpose() * d
+    d = problem.total_constraints()
+        .eval(x, u)
+    return g + v.T * d
         + 0.5 * rho * d.squared_norm()
@@
         dxu = stack(dx[:,i], du[:,i])
-        slope += lqr.q(i).transpose() * dxu
-        slope += dv[:, i].transpose() * lqr.d(i)
-    slope += lqr.q(N).x().transpose() * dx[:, N]
-    slope += dv[:, N].transpose() * lqr.d(N)
+        slope += lqr.q(i).T * dxu
+        slope += dv[:, i].T * lqr.d(i)
+    slope += lqr.q(N).x().T * dx[:, N]
+    slope += dv[:, N].T * lqr.d(N)
     return slope
@@
     alpha = 1
-    m_o = eval_merit(problem, x, u, v, rho)
+    m_o = eval_merit(
+        problem, x, u, v, rho)
     slope = slope(lqr, dx, d, dv, rho)
@@
             problem, x_n, u_n, v_n, rho)
-        if m_n < m_o + kArmijo * alpha * slope:
+        if m_n < m_o +
+            kArmijo * alpha * slope:
             return alpha
@@
         lqr = build_lqr(problem, x, u)
-        dx, du = lqr.Solve()
-        dv = lqr_dual_solve(lqr, dx, du)
+        dx, du, P, p = lqr.Solve()
+        dv = lqr_dual_solve(P, p, dx)
         if dx == 0 and du == 0:
@@
     return x, u, v
-VII Benchmarks
+XII Benchmarks
+Figure 1: The solution of the quad-pendulum problem.
 Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$).
@@
 Below, you can find a visual description of the problem and solution, as well as solver logs.
-Figure 1: The solution of the quad-pendulum problem.
 Iteration Objective $\|c\|^{2}$ $m_{\rho}^{\prime}$ $\alpha$
@@
 33 10.506933 1.0591413968086272e-08 -0.00010727 1.0
-VIII Conclusion
+XIII Conclusion
 This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
@@
 This is due to the fact that inverting a matrix of dimension $d$ takes $O(d^{3})$ sequential time and $O(\log(d))$ parallel time.
-This framework can easily be extended to also handle constrained problems, ideally using an augmented Lagrangian [5] or interior point [12] method; using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
+This framework can easily be extended to also handle constrained problems.
+Ideally, an augmented Lagrangian [5] method would be used to handle general equality constraints and an interior point [12] method would be used to handle general inequality constraints.
+It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian method.
+Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
 Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton.

### v5

@@
 As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6].
-As our algorithm avoids sequential rollouts of the nonlinear dynamics, it can be combined with [17] to run in $O(\log(N))$ parallel time per iteration (where $N$ is the number of stages), as well as $O(1)$ parallel time per line search iteration.
+The solutions of the LQR-shaped subproblems posed by our algorithm can be be parallelized to run in time logarithmic in the number of stages, states, and controls.
+Moreover, as our method avoids sequential rollouts of the nonlinear dynamics, it can run in $O(1)$ parallel time per line search iteration.
 Therefore, this paper provides a practical, theoretically sound, and highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems.
+An open-source JAX implementation of this algorithm can be found on GitHub ($joaospinto/primal\textunderscore dual\textunderscore ilqr$).
 II Introduction
@@
 Our algorithm explores the sparsity structure of the Newton-KKT systems of unconstrained discrete-time optimal control problems, resulting in only $O(N(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem.
-Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n))$ parallel time complexity, using a minor variation of [18], and that the remaining parts of our method have constant parallel time complexity.
+Moreover, we show that the primal-dual Newton-KKT systems we pose can be solved in $O(\log(N)\log(n)+\log(m))$ parallel time complexity, using a minor variation of [17], and that the remaining parts of our method have constant parallel time complexity.
 Our algorithm improves on earlier direct multiple shooting methods, specifically by guaranteeing global convergence and not impeding local superlinear convergence (even without second order corrections), under the assumptions described in [4].
@@
 $m_{\rho}(x,\lambda)=\mathcal{L}(x,\lambda)+\frac{\rho}{2}\lVert c(x)\rVert^{2}.$ (7)
-Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above. Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
+Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above; moreover, $D(\cdot;\cdot)$ is used to represent the directional derivative operator.
+Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
 $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&0\end{pmatrix})$ (8)
@@
 This method of computing the multipliers also has the advantage of involving fewer computations and being more numerically stable.
-XI Algorithm Overview
-In this section, we provide some pseudocode for the sequential version of the proposed algorithm.
-For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
-def build_lqr(problem, x, u, v):
-    n = problem.num_states()
-    m = problem.num_controls()
-    N = problem.num_stages()
-    for i in range(N):
-        cost_grad = problem.cost(i)
-            .gradient()
-            .eval(x[:,i], u[:,i])
-        dyn_grad = problem.dynamics(i)
-            .gradient()
-            .eval(x[:,i], u[:,i])
-        A_i = dyn_grad.T[:,:n]
-        B_i = dyn_grad.T[:,n:]
-        if i == 0:
-            d_i = problem.initial_state()
-                - x[:,0]
-        else:
-            y = problem.dynamics(i-1)
-                .eval(x[:,i-1], u[:,i-1])
-            d_i =  y - x[:,i]
-        l_i = cost_grad +
-            dyn_grad * costates[:,i+1]
-            - costates[:,i]
-        cost_hess = problem.cost(i)
-            .hessian()
-            .eval(x[:,i], u[:,i])
-        dyn_hess = problem.dynamics(i)
-            .hessians()
-            .eval(x[:,i], u[:,i])
-        Q_i = cost_hess
-            + dyn_hess * costates[:,i+1]
-    l_N = problem.cost(N).gradient()
-        .eval(x[:,N]) - costates[:,N]
-    Q_N = problem.cost(N).hessian()
-        .eval(x[:,N])
-    return LQR(A_*, B_*, d_*, Q_*, l_*)
-def lqr_dual_solve(P, p, dx):
-    Np1 = dx.shape[1]
-    return [P[i] * dx[:, i] + p[i]
-        for i in range(Np1)]
-def merit_rho(d, dv):
-    if d == 0:
-        return 0.01
-    return 2 * dv.norm() / d.norm()
-def eval_merit(problem, x, u, v, rho):
-    g = problem.total_cost().eval(x, u)
-    d = problem.total_constraints()
-        .eval(x, u)
-    return g + v.T * d
-        + 0.5 * rho * d.squared_norm()
-def slope(lqr, dx, du, d, dv, rho):
-    N = lqr.num_stages()
-    slope = -rho * d.squared_norm()
-    for i in range(N):
-        dxu = stack(dx[:,i], du[:,i])
-        slope += lqr.q(i).T * dxu
-        slope += dv[:, i].T * lqr.d(i)
-    slope += lqr.q(N).x().T * dx[:, N]
-    slope += dv[:, N].T * lqr.d(N)
-    return slope
-def line_search(problem, lqr, x, u, v,
-                dx, du, dv, rho):
-    alpha = 1
-    m_o = eval_merit(
-        problem, x, u, v, rho)
-    slope = slope(lqr, dx, d, dv, rho)
-    while alpha > kMinStepSize:
-        x_n = x + alpha * dx
-        u_n = u + alpha * du
-        v_n = v + alpha * dv
-        m_n = eval_merit(
-            problem, x_n, u_n, v_n, rho)
-        if m_n < m_o +
-            kArmijo * alpha * slope:
-            return alpha
-    raise LineSearchFailure()
-def primal_dual_ilqr(problem, states_ws,
-        controls_ws, costates_ws):
-    x = states_ws
-    u = controls_ws
-    v = costates_ws
-    for it in range(kMaxIterations):
-        lqr = build_lqr(problem, x, u)
-        dx, du, P, p = lqr.Solve()
-        dv = lqr_dual_solve(P, p, dx)
-        if dx == 0 and du == 0:
-            if dv == 0:
-                return x, u, v
-            else:
-                v += dv
-                continue
-        rho = merit_rho(lqr.d(), dv)
-        alpha = line_search(
-            problem, lqr, x, u,
-            v, dx, du, dv, rho)
-        x += alpha * dx
-        u += alpha * du
-        v += alpha * dv
-    return x, u, v
-XII Benchmarks
+XI Benchmarks
 Figure 1: The solution of the quad-pendulum problem.
-Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$).
+Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from GitHub ($google/trajax$) and modified to treat constraints as high-penalty costs (using a gain of $100$).
 We also increased the final state cost gains on the positions to $1000$.
@@
 33 10.506933 1.0591413968086272e-08 -0.00010727 1.0
-XIII Conclusion
-This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
-Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
-Moreover, it inherits all the desirable properties that NPSQP has been shown to have, such as not impeding local superlinear convergence, global convergence, and the lack of need for second order corrections.
-None of acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6] combine all of these properties.
-The proposed algorithm has sequential time complexity $O(N(n+m)^{3})$, where $N,n,m$ are respectively the number of stages, states, and controls of the given problem, and parallel time complexity $O(\log(m)+\log(N)\log(n)$.
-Note that a naïf application of the NPSQP algorithm would have sequential time complexity $O(N^{3}(n+m)^{3})$ and parallel time complexity $O(\log(N)+\log(n+m))$ (although requiring a much larger number of workers), as the Newton-KKT systems have dimension $(N+1)(n+m)$.
-This is due to the fact that inverting a matrix of dimension $d$ takes $O(d^{3})$ sequential time and $O(\log(d))$ parallel time.
-This framework can easily be extended to also handle constrained problems.
-Ideally, an augmented Lagrangian [5] method would be used to handle general equality constraints and an interior point [12] method would be used to handle general inequality constraints.
-It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian method.
-Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
-Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton.
-Moreover, as it never performs dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.
+XII Conclusion
+This paper introduces a new algorithm for solving unconstrained discrete-time optimal control problems, called Primal-Dual iLQR.
+This framework can easily be extended to handle constrained problems.
+Ideally, an interior point [12] method would be used to handle general inequality constraints.
+It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian [5] method, although local superlinear convergence would be lost.
+In both cases, the LQR shape of the posed subproblems can be preserved.
+Preserving superlinar local convergence in the presence of general equality constraints would require employing a linear equality-constrained LQR subproblem solver.
+Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method (including adding substantially slower iterative subproblem solves).
+Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton, as it treats both state and control trajectories as free variables in the optimization problem (without incurring extra computational costs); this allows them to be independently seeded, even in dynamically infeasible ways.
+The proposed algorithm is globally convergent and does not impede superlinear local convergence, without requiring second order corrections to be applied.
+Finally, solving the subproblems posed by our method can be parallelized to run in time logarithmic in the number of stages, states, and controls.
+Moreover, as our method never performs nonlinear dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.